OpenStudy (anonymous):

Solve. (3x^2)=33x+24 1) 11±√156 / 2 2) 11≠√89 / 2 3) -11± (3√89) / 2 4) -11± (9√17) / 2

OpenStudy (anonymous):

so would that be number 4?

OpenStudy (anonymous):

\[11\pm 3\sqrt{17}\div2\]

OpenStudy (anonymous):

r u sure u have these options.check again please

OpenStudy (anonymous):

ok one sec

OpenStudy (anonymous):

1) should be 11±√153 / 2

OpenStudy (anonymous):

2) 11±√89 / 2

OpenStudy (anonymous):

\[\sqrt{153\[\neq \sqrt{89}\]

OpenStudy (anonymous):

i think the answers are not correct